OpenStudy (anonymous):

lim ( (x)/(x-1) - [(1) / (lnx)] ) x->1

OpenStudy (anonymous):

i think you have to subtract first

OpenStudy (anonymous):

may you please rewrite :D

OpenStudy (anonymous):

\[\lim_{x\to 1^+}\frac{x\ln(x)-(x-1)}{(x-1)\ln(x)}\]

OpenStudy (anonymous):

\[\lim_{x\to 1^+}\frac{x\ln(x)-x+1}{(x-1)\ln(x)}\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} ( \frac{x}{x-1} - \frac{1}{lnx} )\]

OpenStudy (anonymous):

sorry for the confusion

OpenStudy (anonymous):

it's a L'H problem

OpenStudy (anonymous):

got it subtract first, then probably l'hopital

OpenStudy (anonymous):

so if i plug in 1 for x i get 0/0. so... then what do i do exactly?

OpenStudy (anonymous):

i'm suppose to take use L'H, but it's in a weird looking form for that

OpenStudy (anonymous):

remember this is the same as \[(\lim_{x \rightarrow 1}x/x-1)-(\lim_{x \rightarrow 1}1/lnx)\] \[(\lim_{x \rightarrow 1}1/1)-(\lim_{x \rightarrow 1}0/1/x)\] then we get 1-0=1 answer is one, sorry my first answer was wrong

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

answer is 1/2

OpenStudy (anonymous):

\[\lim{x\to 1^+}\frac{1}{\ln(x)}=\infty\] you cannot take the limits separately like that

OpenStudy (anonymous):

start with \[\lim_{x\to 1^+}\frac{x\ln(x)-x+1}{(x-1)\ln(x)}=\lim_{x\to 1^+}\frac{x\ln(x)-x+1}{x\ln(x)-\ln(x)}\] then use l'hopital take the derivative top and bottom separately

OpenStudy (anonymous):

but \[\lim_{x \rightarrow a}f(x)-g(x)=\lim_{x \rightarrow a}f(x)-\lim_{x \rightarrow a}g(x)\] right?

OpenStudy (anonymous):

derivative of the numerator simplifies to \(\ln(x)\) @anonymoustwo44 yes, but both of those limits are infinity, so you are in the form \(\infty-\infty\) so you have to do more work

OpenStudy (anonymous):

what you wrote is true assuming the limits are finite

OpenStudy (anonymous):

@satellite73 ah ok ok

OpenStudy (anonymous):

btw there is no + on the 1

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i didn't follow what you did :(

OpenStudy (anonymous):

i'm looking at x/x-1

OpenStudy (anonymous):

in any case if you take the derivative top and bottom separately you get \[\frac{\ln(x)}{\ln(x)+1-\frac{1}{x}}\] which is still in the form \(\frac{0}{0}\) so you have to do it one more time

OpenStudy (anonymous):

take the derivative top and bottom again you get \[\frac{\frac{1}{x}}{\frac{1}{x}+\frac{1}{x^2}}\] now replace x by 1 and get \[\frac{1}{2}\]

OpenStudy (anonymous):

it's the minus between the fractions i dont know how to deal with...

OpenStudy (anonymous):

back up to the first step

OpenStudy (anonymous):

\[\lim_{x \rightarrow 1} ( \frac{x}{x-1} - \frac{1}{\ln(x)} )=\lim_{x\to 1}\frac{x\ln(x)-x+1}{(x-1)\ln(x)}=\lim_{x\to 1}\frac{x\ln(x)-x+1}{x\ln(x)-\ln(x)}\] this step is pure algebra, just subtraction

OpenStudy (anonymous):

next step is take the derivative top and bottom you get \[\lim_{x\to 1}\frac{\ln(x)}{\ln(x)+1-\frac{1}{x}}\]

OpenStudy (anonymous):

try replacing x by 1 and you see that you get \(\frac{0}{0}\)

OpenStudy (anonymous):

what's happening in that first step

OpenStudy (anonymous):

this one \[\lim_{x \rightarrow 1} ( \frac{x}{x-1} - \frac{1}{\ln(x)} )=\lim_{x\to 1}\frac{x\ln(x)-x+1}{(x-1)\ln(x)}=\lim_{x\to 1}\frac{x\ln(x)-x+1}{x\ln(x)-\ln(x)}\]??

OpenStudy (anonymous):

yes.. after the first =

OpenStudy (anonymous):

i know it's algebra.. that's my problem :(

OpenStudy (anonymous):

that is algebra \[\frac{a}{b}-\frac{c}{d}=\frac{ad-bc}{bc}\]

OpenStudy (anonymous):

i think you mean over bd

OpenStudy (anonymous):

yes right

OpenStudy (anonymous):

multiply the two denominators together to get \[(x-1)\ln(x)=x\ln(x)-\ln(x)\] for our denominator

OpenStudy (anonymous):

numerator is \[x\ln(x)-(x-1)=x\ln(x)-x+1\]

OpenStudy (anonymous):

ok with ya

OpenStudy (anonymous):

so that gives \[\frac{x\ln(x)-x+1}{x\ln(x)-\ln(x)}\] right ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we want to replace x by 1 but we cannot because we will get \(\frac{0}{0}\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so we use l'hopital's rule and take the derivative of the top and bottom. that will give \[\lim_{x\to 1}\frac{\ln(x)}{\ln(x)+1-\frac{1}{x}}\]

OpenStudy (anonymous):

we again try to replace x by 1 and again we get \(\frac{0}{0}\)

OpenStudy (anonymous):

we did product rule in numerator, right?

OpenStudy (anonymous):

well both i guess

OpenStudy (anonymous):

so we have to use l'hopital one more time this time we get \[\frac{\frac{1}{x}}{\frac{1}{x}+\frac{1}{x^2}}\]

OpenStudy (anonymous):

yes product rule in numerator and denominator to find the derivative of \(x\ln(x)\)

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

then that's 1/2... got it. shesh!

OpenStudy (anonymous):

btw, how do you type out the numbers so fast? it takes me forever to use that thing

OpenStudy (anonymous):

i am not using the equation editor

OpenStudy (anonymous):

what's your secret

OpenStudy (anonymous):

i am typing in latex

OpenStudy (anonymous):

so for example if i want \[\lim_{x\to 1}\frac{\ln(x)}{\ln(x)+1-\frac{1}{x}}\] i type in \lim_{x\to 1}\frac{\ln(x)}{\ln(x)+1-\frac{1}{x}}

OpenStudy (anonymous):

by hand? good grief charlie brown

OpenStudy (anonymous):

i started answering questions here to practice my latex almost a year ago, and know i am getting better at it, but i got hooked on answering qustions you enclose the code by first typing "\["

OpenStudy (anonymous):

and end it with "\]" try it if you like if you want to see the code for anything written, right click on it and select "view latex"

OpenStudy (anonymous):

why would you want to get good at latex, is that a valued skill anywhere?

OpenStudy (anonymous):

you can try for example \int_0^{\infty}e^{-x}dx and see what you get

OpenStudy (anonymous):

no it is only good for type setting

OpenStudy (anonymous):

it is the way one writes math, nothing further

OpenStudy (anonymous):

/[\int_0^{\infty}e^{-x}dx]

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

oh well i'll work on it.. ty for the help!